# Why Larp?! A Synthesis Paper on Live Action Roleplay in Relation to HCI Research and Practice

### v1

@@ 1. Introduction
Live Action Roleplay (larp) contains many elements, such as use of technologies, futuring through first-person perspectives, interaction and activity design solutions and game topics, that make it interesting in an Human-Computer Interaction (HCI) context. While similar phenomenona have been around for a long time, larp in its modern form is a cultural expression and a subculture that has been around since the 1980s, in the beginning mostly focused on fantasy settings (Tychsen et al., 2006). The early focus on fantasy in larp might be a contributing factor to the misconception that larp in its essence is mostly non-technological, but despite being an embodied and physical experience, larp explores technologies and futuring in an abundance of ways, and can offer powerful methods for this. In this paper we identify how larp can be relevant in relation to HCI. Larpers have been early adopters of technology, and larp also has been a way of exploring historical, current and future technologies. Larp can be seen as a type of media, a cultural expression and even a tool for learning (Kurz, 2015). In larp, a group of people takes on roles and engages in embodied co-creative storytelling creating an immersive experience together, often enhanced by clothing, props and immersive venues (Salen and Zimmerman, 2003). Within this basic foundation, the purpose, number of participants, themes, type of venues, duration and so on vary greatly between different larps (Tychsen et al., 2006).
@@
The HCI community has for a long time engaged in games and play, simulations, as well as with embodied interaction, public play, augmented reality, roleplaying, and other related concepts that to a lesser or greater extent connect to larp. During the last 20 years, larp as a phenomenon in itself, and its uses for the HCI research community, has gained increasing attention. In 2019 at the Designing Interactive Systems (DIS) conference, a workshop on larp as an embodied design research method was held by a subgroup of the authors of this paper (Márquez Segura et al., 2019). This sparked a lot of discussions and showed that there was an interest in larp within the HCI community. In 2023, the authors of this paper held a workshop at the premier CHI Conference on Human Factors in Computing Systems on educational larp (Robinson et al., 2023), which cemented the need to continue the discourse of unpacking larp and its relationship to HCI. This paper addresses this need by synthesising previous work in the field through first, a practitioner overview of larp in section 4, followed by a more structured literature review within the ACM digital library in section 5. We map the reviewed literature to broader themes and create classifications which highlight a number of research gaps. Through an extensive review of current practices and use of larp in HCI, the article argues that larp is more than a relevant domain of study. We identify unique opportunities that larps bring to HCI related to its capabilities of supporting sensitising, ideation, and exploring possible futures.
@@ 2. Background
This section serves as a brief introduction to larp, and defines what we see as larp, and larp-related activities. This background is brief, as in later sections there will be a more comprehensive overview of larp within the context of HCI as part of our contribution.
@@ 2.1. History of Larp
The cultural phenomenon of larp as we know it today emerged during the early 1980s, based in tabletop roleplaying, with many of the early larps set in fantasy worlds (Tychsen et al., 2006), and, especially in the Nordic countries, also based in improvisational theater (202, 2023). But there is also a long history of larp-like events before modern times, from ancient Egypt (Stenros et al., 2007), China (Contributor, 2021; abs, 2023), Rome (Rangström, 2023) and more, where people dressed up, found immersive locations and and took on roles. From very early on larp like activities became a way to also playfully investigate technological possibilities; One of the world’s first robots, Da Vinci’s robotic armour from 1495, seems to have been built for a divertisement, a larp-like form of entertainment (Rosheim, 2006), and in 1779 an advanced giant fire breathing dragon was built for a roleplay adventure arranged by the Swedish king  [^1] https://digitaltmuseum.se/021046500625/karusell-vid-drottningholm-1779-erovringen-av-galtarklippan.
Larp as a subculture in its modern form grew during the 1980s, initially heavily leaning on fantasy as a theme (Tychsen et al., 2006), but quickly the variations of this versatile media proliferated and genres such as science fiction and dystopic futuristic themes emerged  [^2] https://lajvhistoria.se/arkiv/, http://www.ifgs.org/history.asp.
@@ 2.2. Larp Today
Today larp is a vibrant subculture. A conservative estimate would be that at least 13 million people participate in larps yearly (Xiong et al., 2023), and larps are arranged in many parts of the world; for instance China, (YU, 2021), the US (Dyvers Hands & individual authors, 2014), Europe  [^3] Zero, https://zero.notonlylarp.com/; Nordic Larp in France, https://nordiclarp.org/tag/france/; Mythodea, https://mythodea.de/en/home-en/ (Back, 2014), South Africa (Robnson, 2017), Brazil (Karlsson, 2014), Japan (Kamm, 2020), and Australia (Tychsen et al., 2006).
Larp is a sub-culture, but has also received attention in the wider society, for instance through museum exhibitions  [^4] The Game of the World, https://www.varldskulturmuseet.se/utstallningar/varldens-spel/ and traditional media (doc, 2023; Clasper, 2017). In the experience industry, big corporations such as Disney have given attention to the design possibilities of larp. (Jahromi, 2022; Hon, 2023). Larp-like methods are also used in different educational contexts, then often refered to as edu-larp  [^5] LajvVerkstaden, https://lajvverkstaden.se/ (Schrier, 2019; Bowman, 2014). Outside HCI there is an abundance of research on the benefits of using edu-larp, for instance how it promotes empathy, identity exploration, problem solving, creativity and cooperation (Maragliano, 2019; Lacanienta, 2022; car, 2015; Mochocki, 2023),
@@ 2.3. What Larp Is and What It Is Not
Larp is a versatile format, making it somewhat hard to pin down a clear definition (Tychsen et al., 2006). There have been several previous definitions formulated within the HCI community and in game design, for instance Salen and Zimmerman, already in their 2003 book Rules of Play - Game Design Fundamentals (Salen and Zimmerman, 2003) provide a section about larp, where the definitions for instance include taking on personas of fictional characters, formal game statistics, narrative backstories, real physical spaces, acting out characters action, game mastering, social interaction and pursuing of narrative threads. Our definition in this paper mirrors contemporary trends in the larp community, and tries to encompass the complexity and variation of larps. There are an abundance of other definitions of larp and role playing games (Harviainen et al., 2018; Zagal and Deterding, 2018). In this section we will formulate the definition of larp used in this paper.
@@
It should be noted right away that there exist many larps that do not fulfil all of these criteria – but together they frame not just the most common forms but also capture key aspects of larp that make them an interesting and unique art form. Role-taking is particularly important. Although physical aspects such as venue, costumes and props can play important roles for the immersion, larp is foremost a mental process, engaging with immersing into the story world and state of mind of a character, through an embodied first-person experience. Participants pretend to be someone else, with other feelings and agencies, in a different situation/fiction than the ordinary world around them.
@@
Larping is typically done within some frames given by organizers, such as world setting and a collection of rules for the experience. Larps are thus designed, and have designers that may or may not also be players. Many kinds of larp exist, some are for a few persons  [^6] Stockholm Scenario Festival, https://scenariofestival.se/ up to several thousand participants  [^7] Mythodea, https://mythodea.de/en/home-en/. They differ in duration from an hour up to several days or even weeks. Venues range from a simple classroom to a huge castle complex or an entire forest, and costumes range from simple t-shirts with prints on, to elaborate armor. There is also great variation in narratives, topics and tonality.
@@
In this article we distinguish between leisure larps and edu-larps. Edu-larps in this context we define as when larp methodology is used for specific learning purposes (Bowman et al., 2023). Further, there are larp-like experiences, where the term larp is (commonly) not used, and often without clear connections to the larp community. Sometimes the word larp is even consciously avoided, for fear of the word being seen as unserious. Larp is also partly entwined with other similar sub-cultures such as megagames, cos-play, murder mysteries, some types of AR-games, pervasive games, reenactment, SCA (Society for Creative Anachronism), and escape rooms. Those can, especially from the outside, look a lot like larp. However, most often these larp-like experiences do not focus on role-taking, and therefore fall outside the scope of this article. This article will mostly focus on larps that are explicitly described as larp, but also to some extent engage with other larp-like experiences where the experience fits the definition, even if the term larp is not used.
@@
There are also other kinds of activities in HCI that can be considered larp adjacent, such as designing for mixed reality performances (Zhou et al., 2019; Benford and Giannachi, 2011; Macaulay et al., 2006; Nam and Nitsche, 2014; Taylor et al., 2017) or embodied spectator experiences (Tekin and Reeves, 2017; Seering et al., 2017; Schnädelbach et al., 2008). While these activities can be considered similar to larp in many ways, they differ primarily on the basis of the lack of the first-person perspective in most cases (Striner et al., 2019). Larp, different from performances, do not include an audience and participants have a high level of agency, as co-creating the story experience.
@@ 2.3.1. Different types of larp designs
There is not one cohesive way a larp is designed, nor for instance a single approach to writing larp scripts. Thematically larps vary greatly. Fantasy may dominate, but there are also larps that are set as sci-fi  [^8] Aurum LARP, https://scifixfantasy.com/events-all/aurum-larp/, steampunk  [^9] Steampunk i Sverige, https://steampunkisverige.wordpress.com/category/lajv/, post-apocalyptic (see Figure 2), Regency  [^10] Fortune & Felicity: A Jane Austen Larp, https://anna905.wixsite.com/austenlarp, musical  [^11] Lovestories by ABBA, https://scenariofestival.se/archive/scenarios-2022/lovestories-by-abba/, Wild West  [^12] The Epic Western Larp, http://www.ultimewestern.com/en/accueil-en/, 1920s  [^13] Terra Incognita, https://lajvhistoria.se/lajv/Terra_Incognita.45, and many others. There are also different creative cross-overs, such as Jane Austen meets Westworld  [^14] The Austen Experience, A Winter’s Ball, https://austen.atropos.se/.
Just as in other forms of rpgs (roleplaying games), larps can be heavily scripted, often referred to as railroaded, or very free, often referred to as sandbox (201, 2014; 202, 2021). Larp holds game-like qualities, as there are a set of rules for each larp. The physical enactment of such rules help uphold the simulation of the game world. Rules are often also used to affect the narrative or the players’ out-of-game knowledge. These types of rules are often referred to as meta-techniques (Stark, 2014).
There are some main types of larp: Gamistic battle larps (american style larps), (Unknown, 2022), Nordic larp, a style which has grown to become an influential with a focus on communal storytelling and emotional experience over rules (Saitta et al., 2014), and Jubensha larp which is a Chinese form of larp that has exploded in popularity during recent years and often played as murder mysteries with added roleplaying and immersive venues (Xiong et al., 2023; Xiong, 2022).
Larps, especially those in the Nordic larp style, can have transformative impacts on players, and is a strong method for engaging and creating emotions and transformation (Bowman and Baird, 2022). Some larps are designed with the explicit purpose of (re)learning, practicing prosocial behaviors, and inspiring social change, for example a non-profit larp company Ursula, which has made larps about enslavement, sustainable development, and feminism  [^15] Förenigen Ursula, https://www.foreningenursula.se/, and Not Only Larp, a company that makes immersive experiences for social change, such as larps focused on critiquing capitalism (see Figure 4).
@@ 3. Method
The methodology of this work, especially the synthesising section, can be characterized in the tradition of Humanistic HCI (Bardzell and Bardzell, 2016). A group of experts engaged in the sensemaking and contextualization for this piece together, in an iterative process. The expert group consists of HCI researchers from the US and Europe, in different stages of their academic career, all who have published larp related research. Several of the authors are also active larpers and larp designers, with a long history of studying and arranging larps, within and outside the HCI context. The expert group of researchers consisted of: 3 professors, 2 assistant professors, 1 postdoc and 2 PhD students. All members of the expert group are co-authors of this paper. There is a limited amount of work done on larp in relation to HCI, and combined the expert group has been involved in a majority of the research in this field that has been done. During the span of several months, from fall 2022 until august 2023, this expert group met around once a month and brought together knowledge on larp, and technology in relation to larp, through online meetings and conversations. At CHI’23 the group arranged the workshop called edu-larp@CHI, focused on educational larps, bringing together 15 researchers with interest on the topic. This workshop further informed and spurred the discussions in the expert group.
The expert group brought together and compared their knowledge on larp and larp related research, spanning several decades, and also presented and discussed larp research exemplars, as well as together identified research interests and topics relevant in relation to larp and HCI.
Through these discussions the group identified knowledge on larp to synthesise, toward creating a foundation for future research in the HCI community. This included a literature review on larp-related work published within ACM, as well as a synthesis of knowledge from larp practice outside academia, and a set of annotated larp research exemplars, as presented in this paper.
@@ 3.1. Positionality Statement
The authors are, in most cases, insiders in different larp communities. This allows for deeper insights into larp, however we also acknowledge that this can create a bias toward our Western-centric perspective of larp. Further, being the designers and larpers the authors have preferred methods, genres and larp styles, this especially affects the research design exemplars. The co-authors of this paper are located in North America and Europe, and have only directly experienced larp as it is created and played in these regions. While we don’t have first hand experience with larp outside of NA and Europe, we know there is also robust Chinese and Brazilian larp communities, which we do make sure to include in our practitioner overview in Section 4. The focus on Europe and the US limits the perspective of this synthesis paper to these traditions, apart from citing work that has been published in the HCI research literature. The authors’ long term engagements with specific larp communities, as well as with larp related HCI research, has shaped their view on larp, and despite trying to provide an overview, we acknowledge that this overview is not objective or entirely disconnected from the positionalities of the authors.

### v2

@@ 1. Introduction
Live Action Roleplay (larp) contains many elements that make it interesting in the context of Human-Computer Interaction (HCI), such as use of technologies, futuring through first-person perspectives, interaction and activity design solutions and game topics. While similar phenomenon have been around for a long time, larp in its modern form is a cultural expression and a subculture that has been around since the 1980s. In the beginnings, it mostly focused on fantasy settings (Tychsen et al., 2006). The early focus on fantasy in larp might be a contributing factor to the misconception that larp in its essence is mostly non-technological, but despite being an embodied and physical experience, larp explores technologies and futuring in an abundance of ways, and can offer powerful methods for this. In this paper we identify how larp can be relevant in relation to HCI. Larpers have been early adopters of technology, and larp also has been a way of exploring historical, current and future technologies. Larp can be seen as a type of media, a cultural expression and even a tool for learning (Kurz, 2015). In larp, a group of people take on roles and engage in embodied co-creative storytelling creating an immersive experience together, often enhanced by clothing, props and immersive venues (Salen and Zimmerman, 2003). Within this basic foundation, the purpose, number of participants, themes, type of venues, duration and so on vary greatly between different larps (Tychsen et al., 2006).
@@
For many years, the HCI community has engaged with topics such as games and play, simulations, embodied interaction, public play, augmented reality, roleplaying, and other related concepts that to a lesser or greater extent connect to larp. During the last 20 years, larp as a phenomenon in itself, and its uses for the HCI research community, has gained increasing attention. In 2019 at the Designing Interactive Systems (DIS) conference, a workshop on larp as an embodied design research method was held by a subgroup of the authors of this paper (Márquez Segura et al., 2019). This sparked a lot of discussions and showed that there was an interest in larp within the HCI community. In 2023, the authors of this paper held a workshop at the premier CHI Conference on Human Factors in Computing Systems on educational larp (Robinson et al., 2023), which cemented the need to continue the discourse of unpacking larp and its relationship to HCI. This paper addresses this need by synthesising previous work in the field through first, a practitioner overview of larp in section 4, followed by a more structured literature review within the ACM digital library in section 5. Next, in section 6, we present four larp research exemplars of our own work which show a broad range of how larp can be used in HCI. Lastly, we map the reviewed literature to broader themes and create classifications which highlight a number of research gaps in section 7. Through an extensive review of current practices and use of larp in HCI, the article argues that larp is more than a relevant domain of study. We identify unique opportunities that larps bring to HCI related to its capabilities of supporting sensitising, ideation, and exploring possible futures.
@@ 2. Background
This section serves as a brief introduction to larp, defining what we see as larp and larp-related activities. This background is brief, as in later sections there will be a more comprehensive overview of larp within the context of HCI as part of our contribution.
@@ 2.1. History of Larp
The cultural phenomenon of larp as we know it today emerged during the early 1980s, based in tabletop roleplaying, with many of the early larps set in fantasy worlds (Tychsen et al., 2006), and, especially in the Nordic countries, also based in improvisational theater (202, 2023). But there is also a long history of larp-like events before modern times, from ancient Egypt (Stenros et al., 2007), China (Contributor, 2021; abs, 2023), Rome (Rangström, 2023) and more, where people dressed up, found immersive locations and took on roles. From very early on, larp-like activities became a way to also playfully investigate technological possibilities; One of the world’s first robots, Da Vinci’s robotic armour from 1495, seems to have been built for a divertissement, a larp-like form of entertainment (Rosheim, 2006), and in 1779 an advanced giant fire breathing dragon was built for a roleplay adventure arranged by the Swedish king  [^1] Karusell vid Drottningholm 1779, https://digitaltmuseum.se/021046500625/karusell-vid-drottningholm-1779-erovringen-av-galtarklippan.
Larp as a subculture in its modern form grew during the 1980s, initially heavily leaning on fantasy as a theme (Tychsen et al., 2006), but quickly the variations of this versatile media proliferated and genres such as science fiction and dystopic futuristic themes emerged  [^2] History of the IFGS http://www.ifgs.org/history.asp.
@@ 2.2. Larp Today
Today, larp is a vibrant subculture. A conservative estimate would be that at least 13 million people participate in larps yearly (Xiong et al., 2023), and larps are arranged in many parts of the world; for instance China, (YU, 2021), the USA (Dyvers Hands & individual authors, 2014), Europe such as Zero, (see 23 in Appendix A); Nordic Larp in France  [^3] https://nordiclarp.org/tag/france/; Mythodea, (see 15 in Appendix A) (Back, 2014), South Africa (Robnson, 2017), Brazil (Karlsson, 2014), Japan (Kamm, 2020), and Australia (Tychsen et al., 2006).
Larp is a sub-culture, but has also received attention in the wider society, through museum exhibitions  [^4] The Game of the World, https://www.varldskulturmuseet.se/utstallningar/varldens-spel/ and traditional media (doc, 2023; Clasper, 2017). In the experience industry, big corporations such as Disney have given attention to the design possibilities of larp (Jahromi, 2022; Hon, 2023). Larp-like methods are also used in different educational contexts, then often referred to as edu-larp  [^5] LajvVerkstaden, https://lajvverkstaden.se/ (Schrier, 2019; Bowman, 2014). Outside HCI there is an abundance of research on the benefits of using edu-larp, for instance how it promotes empathy, identity exploration, problem solving, creativity and cooperation (Maragliano, 2019; Lacanienta, 2022; car, 2015; Mochocki, 2023),
@@ 2.3. What Larp Is and What It Is Not
Larp is a versatile format, making it somewhat hard to pin down a clear definition (Tychsen et al., 2006). There have been several previous definitions formulated within the HCI community and in game design, for instance Salen and Zimmerman, already in their 2003 book Rules of Play - Game Design Fundamentals (Salen and Zimmerman, 2003) provide a section about larp, where the definition of larp includes taking on personas of fictional characters, formal game statistics, narrative backstories, real physical spaces, acting out characters action, game mastering, social interaction and pursuing of narrative threads. Our definition in this paper mirrors contemporary trends in the larp community, and tries to encompass the complexity and variation of larps. There are an abundance of other definitions of larp and role playing games (Harviainen et al., 2018; Zagal and Deterding, 2018). In this section we will formulate the definition of larp used in this paper.
@@
It should be noted right away that there exist many larps that do not fulfil all of these criteria – but together they frame not just the most common forms but also capture key aspects of larp that make them an interesting and unique art form. Role-taking is particularly important. Although physical aspects such as venue, costumes and props can play important roles for the immersion, larp is first and foremost a mental process, engaging with immersing into the story world and state of mind of a character, through an embodied first-person experience. Participants pretend to be someone else, with other feelings and agencies, in a different situation/fiction than the ordinary world around them.
@@
Larping is typically done within some frames given by organizers, such as world setting and a collection of rules for the experience. Larps are thus designed, and have designers that may or may not also be players. Many kinds of larps exist, some are for a few persons such as the Stockholm Scenario Festival  [^6] Stockholm Scenario Festival, https://scenariofestival.se/ and others up to several thousand participants such as Mythodea (see 15 in Appendix A). Larps can differ in duration from an hour up to several days or even weeks. Venues range from a simple classroom to a huge castle complex or an entire forest, and costumes range from simple t-shirts with prints on, to elaborate armor. There is also great variation in narratives, topics and tonality.
@@
In this article we distinguish between leisure larps and edu-larps. Edu-larps in this context we define as when larp methodology is used for specific learning purposes (Bowman et al., 2023). Further, there are larp-like experiences, where the term larp is (commonly) not used, and often without clear connections to the larp community. Sometimes the word larp is even consciously avoided, for fear of the word being seen as not serious. Larp is also partly entwined with other similar sub-cultures such as megagames, cos-play, murder mysteries, some types of AR-games, pervasive games, reenactment, SCA (Society for Creative Anachronism)  [^7] Society for Creative Anachronism, https://www.sca.org/, and escape rooms. Those can, especially from the outside, look a lot like larp. However, most often these larp-like experiences do not focus on role-taking, and therefore fall outside the scope of this article. This article will mostly focus on larps that are explicitly described as larp, but also to some extent engage with other larp-like experiences where the experience fits the definition, even if the term larp is not used.
@@
There are also other kinds of activities in HCI that can be considered larp adjacent, such as designing for mixed reality performances (Zhou et al., 2019; Benford and Giannachi, 2011; Macaulay et al., 2006; Nam and Nitsche, 2014; Taylor et al., 2017) or embodied spectator experiences (Tekin and Reeves, 2017; Seering et al., 2017; Schnädelbach et al., 2008). Simulations are also a genre that can be considered adjacent to larp—drawing on roleplaying for applied domains such as within medical fields or for disaster response teams (Toups Dugas et al., 2016; Bowman, 2014). While these activities can be considered similar to larp in many ways, they differ primarily on the basis of the lack of the first-person perspective in most cases (Striner et al., 2019). Larp, different from performances, do not include an audience and participants have a high level of agency, as they are co-creating the story experience.
@@ 2.3.1. Different types of larp designs
There is not one cohesive way a larp is designed, nor a single approach to writing larp scripts. Thematically, larps can vary greatly. The fantasy genre may dominate, but there are also larps that are set as sci-fi (see 4 in Appendix A), steampunk  [^8] Steampunk i Sverige, https://steampunkisverige.wordpress.com/category/lajv/, post-apocalyptic (see Figure 2), Regency (see 9 in Appendix A), musical (see 13 in Appendix A), Wild West (see 20 in Appendix A), 1920s (see 18 in Appendix A), and many others. There are also different creative cross-overs, such as Jane Austen meets Westworld (see 19 in Appendix A).
Just as in other forms of rpgs (roleplaying games), larps can be heavily scripted, often referred to as railroaded, or very free-form, often referred to as sandbox (201, 2014; 202, 2021). Larp holds game-like qualities, as there are a set of rules for each larp. The physical enactment of such rules help uphold the simulation of the game world. Rules are often also used to affect the narrative or the players’ out-of-game knowledge. These types of rules are often referred to as meta-techniques (Stark, 2014).
There are some main types of larp: Gamistic battle larps (american style larps) (Unknown, 2022), Nordic larp, a style which has grown to become an influential with a focus on communal storytelling and emotional experience over rules (Saitta et al., 2014), and Jubensha larp which is a Chinese form of larp that has exploded in popularity during recent years and often played as murder mysteries with added roleplaying and immersive venues (Xiong et al., 2023; Xiong, 2022).
Larps, especially those in the Nordic larp style, can have transformative impacts on players, and are a strong method for engaging and creating emotions and transformation (Bowman and Baird, 2022). Some larps are designed with the explicit purpose of (re)learning, practicing prosocial behaviors, and inspiring social change. For example, a non-profit larp company called Ursula has made larps about enslavement, sustainable development, and feminism  [^9] Förenigen Ursula, https://www.foreningenursula.se/, and Not Only Larp is a company that makes immersive experiences for social change, such as larps focused on critiquing capitalism (see 23 in Appendix A).
@@ 3. Method
The methodology of this work, especially the synthesising section, can be characterized in the tradition of Humanistic HCI (Bardzell and Bardzell, 2016). Humanistic HCI, as with other well established humanistic research methods such as ethno-methodology (O’Reilly, 2012), auto-ethnography (Marechal, 2010; Wall, 2006), and narrative inquiry (Riessman, 1993), this approach embraces the subjectivity of studying human interactions and social phenomenons, thus not striving for objectivity.
A group of experts engaged in the sensemaking and contextualization for this piece together, in an iterative process. During the span of several months, from October 2022 until August 2023, this expert group met around once a month and brought together knowledge on larp, and technology in relation to larp, through online meetings and conversations. At CHI’23 the group arranged the workshop called edu-larp@CHI, focused on educational larps, bringing together 15 researchers with interest on the topic. This workshop, although not part of the data gathering for this paper, further spurred the discussions in the expert group.
Over several months the expert group brought together and compared their knowledge on larp and larp related research, spanning several decades, and also presented and discussed larp research exemplars, as well as together identified research interests and topics relevant in relation to larp and HCI. Through these discussions the group identified knowledge on larp to synthesise, toward creating a foundation for future research in the HCI community. This included a literature review on larp-related work published within ACM, as well as a synthesis of knowledge from larp practice outside academia, and a set of annotated larp research exemplars, as presented in this paper.
@@ 3.1. Positionality Statement
The expert group consists of HCI researchers from the US and Europe, in different stages of their academic career, all who have published larp related research. Several of the authors are also active larpers and larp designers, with a long history of studying and arranging larps, within and outside the HCI context. At the time of writing, the expert group of researchers consisted of: 2 full professors (Sweden and the US), 1 associate professor (Sweden), (2 assistant professors (1 in Sweden, one in Spain), 1 postdoc (the US) and 2 PhD students (in the US and Sweden). All members of the expert group are co-authors of this paper. There is a limited amount of work done on larp in relation to HCI, and combined the expert group has been involved in a majority of the research in this field that has been done.
The authors are, in most cases, insiders in different larp communities. This allows for deeper insights into larp, however we also acknowledge that this can create a bias toward our Western-centric perspectives of larp. Further, being the designers and larpers the authors have preferred methods, genres and larp styles, this especially affects the research design exemplars. The co-authors of this paper are located in North America and Europe, and have only directly experienced larp as it is created and played in these regions. While we don’t have first hand experience with larp outside of NA and Europe, we know there is also robust Chinese and Brazilian larp communities, which we do make sure to include in our practitioner overview in Section 4. The focus on Europe and the US limits the perspective of this synthesis paper to these traditions, apart from citing work that has been published in the HCI research literature. The authors’ long term engagements with specific larp communities, as well as with larp related HCI research, has shaped their view on larp, and despite trying to provide an overview, we acknowledge that this overview is not objective or entirely disconnected from the positionalities of the authors.